OpenStudy (anonymous):

The local zoo has two water tanks for the elephant exhibit that are leaking One water tank contains 12 gal of water and is leaking at a constant rate of 3 gal/h. The second water tank contains 8 gal of water and is leaking at a constant rate of 5 gal/h. When will the two tanks have the same amount of water? Explain. Let x = the number of hours the tanks are filling and let y = the number of gallons in the tank.

OpenStudy (anonymous):

A(h)=(rate)(number of hours)+Starting amount of water \[r_1=-3\] \[r_2=-5\] \[y_1=12\] \[y_2=8\] \[A(t)=-3h+12\] \[B(t)=-5h+8\]

OpenStudy (anonymous):

to find out when they equal the same amount, create an equality -3h+12=-5h+8 \[2h=-4\] \[h=-2\] these two water holding devices will never have the same amount

OpenStudy (anonymous):

thanks man appericate it needed that one